a chemist wants to mix 20% acid solution with a 44% acid solution to get 24 L of a 29% acid solution. How many liters of the 20% solution and how many liters of the 44% solution should be mixed?
Answer by stanbon(75887)

a chemist wants to mix 20% acid solution with a 44% acid solution to get 24 L of a 29% acid solution. How many liters of the 20% solution and how many liters of the 44% solution should be mixed?
---------------------
Quantity Equation: x + y = 24 L
Acid Equation: 0.20x + 0.44y = 0.29*24
---------------------------------------
Rearrange and Modify:
20x + 20y = 20*24
20x + 44y = 29*24
---------------------
Subtract 1st from 2nd to solve for "y":
24y = 9*24
y = 9 L (Amt. of 44% solution in the mixture)
x = 24-9 = 15 L (Amt. of 20% solution in the mixture)
===================
Cheers,
Stan H.